the length of a football field is 7 yards longer than double the width.if the perimeter of the rectangular playing field is 326 yards what are the demensions?
Answer by checkley77(12844)

L=2W+7
2L+2W=326
2(2W+7)+2W=326
4W+14+2W=326
6W=326-14
6W=312
W=312/6
W=52 ANS. FOR THE WIDTH.
L=2*52+7=104+7=111 ANS. FOR THE LENGTH.
PROOF:
2*111+2*52=326
222+104=326
326=326